At noon, a car B is 10 km north of another car A. Car A travels to the east at
30 km/h and car B travels south at 40 km/h. When will they be the closest to each other?

Answer by mananth(16949)

The shortest distance between them will be when Car B reaches the starting point of car A
Car B d=10
r=40km/h
t=d/r
10/40=1/4 hour
at 12:15 pm
m.ananth@hotmail.ca